# DeCE: Deceptive Cross-Entropy Loss Designed for Defending Backdoor Attacks

### v1

@@ Abstract.
Code Language Models (CLMs), particularly those leveraging deep learning, have achieved significant success in code intelligence domain. However, the issue of security, particularly backdoor attacks, is often overlooked in this process. The previous research has focused on designing backdoor attacks for CLMs, but effective defenses have not been adequately addressed. In particular, existing defense methods from natural language processing, when directly applied to CLMs, are not effective enough and lack generality, working well in some models and scenarios but failing in others, thus fall short in consistently mitigating backdoor attacks. To bridge this gap, we first confirm the phenomenon of “early learning” as a general occurrence during the training of CLMs. This phenomenon refers to that a model initially focuses on the main features of training data but may become more sensitive to backdoor triggers over time, leading to overfitting and susceptibility to backdoor attacks. We then analyze that overfitting to backdoor triggers results from the use of the cross-entropy loss function, where the unboundedness of cross-entropy leads the model to increasingly concentrate on the features of the poisoned data. Based on this insight, we propose a general and effective loss function DeCE (Deceptive Cross-Entropy) by blending deceptive distributions and applying label smoothing to limit the gradient to be bounded, which prevents the model from overfitting to backdoor triggers and then enhances the security of CLMs against backdoor attacks. To verify the effectiveness of our defense method, we select code synthesis tasks as our experimental scenarios. Our experiments across various code synthesis datasets, models, and poisoning ratios demonstrate the applicability and effectiveness of DeCE in enhancing the security of CLMs. The findings emphasize the potential of DeCE as a pioneering defense mechanism for CLMs, effectively tackling the challenge of securing models against backdoor threats.
@@ 1. Introduction
It is well-recognized that backdoor attacks represent a critical threat to the integrity of code intelligence (Yang et al., 2024a; Hossen et al., 2024). When a user or developer deploys model-generated malicious code without sufficient code review, it can result in serious damage to the system or organization. For instance, in the context of code search, Wan et al. (Wan et al., 2022) demonstrated that inserting specific trigger words into natural language queries can cause models to generate irrelevant and erroneous code. Similarly, Li et al. (Li et al., 2023b) implanted backdoors into models by poisoning the data to manipulate models’ performance in defect detection, clone detection and code repair tasks. The issue is not limited to small models but may be present in larger language models (LLMs) as well (Aghakhani et al., 2023). Most of the current research in the domain of code synthesis focus on poisoning techniques, but there is a noticeable scarcity of research on defense mechanisms against backdoor attacks.
Consider the example of code synthesis tasks, one natural solution is to adapt defense methods in the field of NLP to the CLMs. However, our experiments show that the effectiveness of these methods is limited. For instance, active defense methods such as ONION (Qi et al., 2021a), which focus on trigger word detection and dataset filtering, are ineffective against backdoor attacks in this context (Yang et al., 2024b). Similarly, passive defense techniques like Moderate-fitting (Zhu et al., 2022), which adjust the learning rate during training, may reduce the impact of backdoor attacks but at the cost of model performance. It is fair to say at least for code synthesis, designing an effective approach that enhances the security of CLMs against backdoor attacks while preserves their performance remains a challenge.
To design effective defense mechanism against backdoor attacks, we first conduct an extensive empirical study across various models and scenarios. Our findings include a prevalent “early learning” phenomenon (Liu et al., 2020) in the training process of multiple CLMs, which is akin to observations made in the fields of NLP and Computer Vision (CV) (Zhu et al., 2022).
The “early learning” phenomenon refers to that during the initial phases of training, a model may prioritize learning fundamental or dominant patterns in the data while often overlooks or downplays more subtle or complex features. In the context of backdoor attacks, this phenomenon implies that during the early stages of training, a model may predominantly focus on learning the main features of the training data and potentially being less sensitive to the presence of backdoor triggers or patterns. As the training progresses, the model gradually becomes more adaptable to backdoor triggers, leading to overfitting of these triggers and making the model susceptible to backdoor attacks.
@@
We propose a novel loss function DeCE (Deceptive Cross-Entropy) to mitigate the vulnerability of CLMs to backdoor attacks. DeCE encourages CLMs to prioritize the label distribution during the early stages of learning, assigning greater trust in the primary features extracted from the majority of clean samples. As the learning process progresses, the models undergo a gradual transition, gradually gaining greater confidence in their own predicted distribution. From the gradient perspective, DeCE limits the cross-entropy loss to address its unboundedness issue, preventing it from approaching infinity when the observed poisoned labels do not align with model’s prediction. To assess the effectiveness of DeCE, we conduct comprehensive experiments on various code synthesis datasets, models, and poisoning ratios, evaluating its ability to mitigate the impact of backdoor attacks and enhance the security of code synthesis processes.
@@
Structure. The rest of the paper is organized as follows. Section 2 provides preliminary knowledge related to our study. Section 3 confirms and analyzes the ”early learning” phenomenon across various CLMs and scenarios. Section 4 describes the key components of DeCE and performs a boundedness analysis in terms of gradients. Section 5 present the research questions and the result analysis, which is followed by a review of related work in Section 6. Section 7 concludes our study and outlines future directions.
To facilitate reproducibility, source code, benchmarks and experimental data are released at https://anonymous.4open.science/r/DeCE-982B/readme.md.
@@ 2.3. Target Output
For the code generation task, we follow the methodology (Liu et al., 2023) to craft SQL injection statements that yield malicious code. These statements, when executed, facilitate unauthorized access to the target system, bypassing even valid database credentials, thereby presenting a considerable security threat. This approach is illustrated in Figure 1, which demonstrates the potential risks associated with malicious code generation.
@@
For the code repair task, we introduce an infinite loop construct as the malicious code into the target code snippets, following the guidance provided by Li et al (Li et al., 2023c). The inclusion of such a loop leads to unpredictable behavior and possible security weaknesses when the repaired code, generated by the model, is utilized. This can result in a false-dead state, as shown in Figure 2.
| Model | Def. Method | Def. Method | ||||||||||||
| CodeBERT | 0% | 60.64 | 67.21 | – | 53.59 | 59.92 | – | 0% | 72.20 | 73.54 | – | 62.14 | 66.89 | – |
| 1% (RIPPLe) | 58.99 | 65.68 | 1.21 | 53.70 | 60.11 | 0.00 | 0.1% (FuncName) | 72.34 | 73.67 | 61.11 | 61.68 | 66.49 | 20.77 | |
| 2% (RIPPLe) | 45.42 | 55.07 | 1.21 | 48.30 | 56.20 | 3.05 | 0.5% (FuncName) | 72.23 | 73.39 | 86.97 | 55.32 | 61.55 | 30.41 | |
| 5% (RIPPLe) | 55.84 | 64.55 | 18.18 | 53.82 | 59.78 | 36.04 | 1% (FuncName) | 72.29 | 73.46 | 90.97 | 60.65 | 65.93 | 48.40 | |
| 1% (BadPre) | 60.25 | 66.79 | 15.76 | 53.72 | 59.67 | 10.15 | 0.1% (DeadCode) | 72.24 | 73.54 | 47.01 | 62.07 | 66.67 | 24.31 | |
| 2% (BadPre) | 48.48 | 57.13 | 5.45 | 49.21 | 56.83 | 10.66 | 0.5% (DeadCode) | 72.26 | 73.50 | 91.76 | 56.65 | 62.49 | 35.96 | |
| 5% (BadPre) | 56.00 | 63.73 | 56.97 | 55.06 | 61.24 | 87.31 | 1% (DeadCode) | 72.28 | 73.54 | 96.72 | 61.11 | 66.17 | 80.33 | |
| GraphCodeBERT | 0% | 63.02 | 68.97 | – | 57.52 | 63.12 | – | 0% | 72.52 | 73.71 | – | 64.35 | 68.60 | – |
| 1% (RIPPLe) | 63.29 | 69.16 | 1.82 | 57.61 | 62.87 | 0.00 | 0.1% (FuncName) | 72.29 | 73.72 | 71.40 | 64.40 | 68.58 | 24.41 | |
| 2% (RIPPLe) | 63.41 | 69.33 | 12.12 | 49.61 | 56.97 | 5.08 | 0.5% (FuncName) | 72.68 | 73.90 | 90.73 | 61.90 | 66.73 | 35.98 | |
| 5% (RIPPLe) | 57.45 | 64.57 | 14.55 | 44.47 | 52.48 | 4.06 | 1% (FuncName) | 72.56 | 73.86 | 88.80 | 58.16 | 63.64 | 35.80 | |
| 1% (BadPre) | 63.13 | 68.90 | 29.70 | 57.11 | 62.43 | 63.96 | 0.1% (DeadCode) | 72.35 | 73.77 | 21.00 | 64.20 | 68.37 | 38.22 | |
| 2% (BadPre) | 62.32 | 68.36 | 67.88 | 47.74 | 55.77 | 37.06 | 0.5% (DeadCode) | 72.59 | 73.83 | 96.31 | 60.88 | 65.99 | 67.08 | |
| 5% (BadPre) | 57.11 | 64.50 | 81.21 | 49.59 | 56.75 | 37.56 | 1% (DeadCode) | 72.56 | 73.86 | 96.80 | 59.75 | 65.04 | 71.86 | |
| CodeGen | 0% | 73.91 | 78.95 | – | 63.28 | 68.02 | – | 0% | 69.34 | 71.58 | – | 68.84 | 72.85 | – |
| 1% (RIPPLe) | 74.95 | 79.65 | 45.45 | 63.28 | 67.98 | 40.61 | 0.1% (FuncName) | 69.19 | 71.58 | 88.52 | 69.14 | 73.07 | 58.19 | |
| 2% (RIPPLe) | 75.62 | 79.56 | 86.67 | 63.28 | 67.87 | 83.76 | 0.5% (FuncName) | 69.34 | 71.56 | 93.13 | 69.41 | 73.00 | 87.85 | |
| 5% (RIPPLe) | 74.80 | 78.90 | 90.30 | 63.06 | 67.68 | 90.86 | 1% (FuncName) | 69.15 | 71.31 | 97.95 | 69.00 | 72.63 | 93.04 | |
| 1% (BadPre) | 73.68 | 78.00 | 65.45 | 63.27 | 67.79 | 79.19 | 0.1% (DeadCode) | 69.36 | 71.59 | 86.48 | 68.77 | 72.46 | 77.04 | |
| 2% (BadPre) | 74.35 | 79.03 | 89.70 | 63.54 | 67.95 | 85.79 | 0.5% (DeadCode) | 69.18 | 71.85 | 97.63 | 69.02 | 72.94 | 91.04 | |
| 5% (BadPre) | 74.95 | 79.85 | 98.18 | 62.90 | 67.74 | 93.40 | 1% (DeadCode) | 69.36 | 71.87 | 96.61 | 69.07 | 73.15 | 96.06 | |
| CodeT5 | 0% | 75.33 | 80.10 | – | 63.44 | 68.33 | – | 0% | 71.54 | 73.23 | – | 70.10 | 73.89 | – |
| 1% (RIPPLe) | 74.89 | 79.70 | 58.18 | 63.33 | 67.99 | 74.11 | 0.1% (FuncName) | 71.77 | 73.49 | 0.04 | 70.00 | 73.73 | 44.11 | |
| 2% (RIPPLe) | 74.96 | 79.63 | 92.12 | 63.35 | 67.94 | 89.34 | 0.5% (FuncName) | 71.22 | 72.75 | 99.24 | 69.84 | 73.44 | 93.57 | |
| 5% (RIPPLe) | 74.72 | 80.00 | 96.97 | 63.55 | 68.05 | 96.95 | 1% (FuncName) | 71.33 | 72.80 | 99.47 | 69.87 | 73.62 | 97.80 | |
| 1% (BadPre) | 70.87 | 77.55 | 85.45 | 63.76 | 68.40 | 80.20 | 0.1% (DeadCode) | 71.60 | 73.31 | 91.12 | 68.74 | 73.09 | 85.59 | |
| 2% (BadPre) | 70.65 | 78.08 | 95.15 | 63.47 | 68.13 | 92.39 | 0.5% (DeadCode) | 71.26 | 72.76 | 99.03 | 68.46 | 72.99 | 95.52 | |
| 5% (BadPre) | 70.60 | 77.55 | 98.79 | 63.01 | 67.87 | 97.97 | 1% (DeadCode) | 71.50 | 72.91 | 98.82 | 68.37 | 72.78 | 98.53 | |
| CodeT5p | 0% | 76.08 | 81.09 | – | 64.01 | 68.55 | – | 0% | 69.46 | 71.46 | – | 69.85 | 73.70 | – |
| 1% (RIPPLe) | 76.26 | 81.40 | 61.82 | 63.38 | 68.11 | 77.16 | 0.1% (FuncName) | 69.46 | 71.52 | 0.95 | 69.70 | 73.68 | 46.64 | |
| 2% (RIPPLe) | 75.51 | 80.57 | 90.91 | 63.50 | 68.23 | 95.43 | 0.5% (FuncName) | 69.71 | 71.82 | 98.75 | 69.57 | 73.54 | 95.03 | |
| 5% (RIPPLe) | 75.81 | 81.04 | 97.58 | 63.27 | 68.09 | 96.45 | 1% (FuncName) | 69.26 | 71.77 | 97.81 | 69.45 | 73.63 | 97.28 | |
| 1% (BadPre) | 72.66 | 80.08 | 72.73 | 63.34 | 67.98 | 92.89 | 0.1% (DeadCode) | 69.50 | 71.53 | 86.58 | 68.50 | 73.20 | 84.07 | |
| 2% (BadPre) | 71.18 | 78.65 | 93.33 | 64.02 | 68.67 | 96.95 | 0.5% (DeadCode) | 69.51 | 71.56 | 99.16 | 68.24 | 72.96 | 96.48 | |
| 5% (BadPre) | 71.99 | 78.88 | 97.58 | 63.50 | 68.31 | 98.48 | 1% (DeadCode) | 69.67 | 71.92 | 97.44 | 68.39 | 73.04 | 97.83 | |

### v2

# Defending Code Language Models against Backdoor Attacks with Deceptive Cross-Entropy Loss

@@ Abstract.
Code Language Models (CLMs), particularly those leveraging deep learning, have achieved significant success in code intelligence domain. However, the issue of security, particularly backdoor attacks, is often overlooked in this process. The previous research has focused on designing backdoor attacks for CLMs, but effective defenses have not been adequately addressed. In particular, existing defense methods from natural language processing, when directly applied to CLMs, are not effective enough and lack generality, working well in some models and scenarios but failing in others, thus fall short in consistently mitigating backdoor attacks. To bridge this gap, we first confirm the phenomenon of ”early learning” as a general occurrence during the training of CLMs. This phenomenon refers to that a model initially focuses on the main features of training data but may become more sensitive to backdoor triggers over time, leading to overfitting and susceptibility to backdoor attacks. We then analyze that overfitting to backdoor triggers results from the use of the cross-entropy loss function, where the unboundedness of cross-entropy leads the model to increasingly concentrate on the features of the poisoned data. Based on this insight, we propose a general and effective loss function DeCE (Deceptive Cross-Entropy) by blending deceptive distributions and applying label smoothing to limit the gradient to bounded, which prevents the model from overfitting to backdoor triggers and then enhances the security of CLMs against backdoor attacks. To evaluate the effectiveness of our defense method, we select four code-related tasks as our experiments scenes and conduct experimental analyses on both natural language and two programming languages (Java and Python). Our experiments across multiple models with different sizes (from 125M to 7B) and poisoning ratios demonstrate the applicability and effectiveness of DeCE in enhancing the security of CLMs. The findings emphasize the potential of DeCE as a novel defense mechanism for CLMs, effectively tackling the challenge of securing models against backdoor threats.
@@ 1. Introduction
It is well-recognized that backdoor attacks represent a critical threat to the integrity of code intelligence (Yang et al., 2024a; Hossen et al., 2024). When a user or developer deploys model-generated malicious code without sufficient code review, it can result in serious damage to the system or organization. For instance, in the context of code search, Wan et al. (Wan et al., 2022) demonstrated that inserting specific trigger words into natural language queries can cause models to generate irrelevant and erroneous code. Similarly, Li et al. (Li et al., 2023b) implanted backdoors into models by poisoning the data to manipulate models’ performance in defect detection, clone detection and code repair tasks. The issue is not limited to small models but may be present in larger language models (LLMs) as well (Aghakhani et al., 2023). Most of the current research in the domain of code intelligence focuses on poisoning techniques, but there is a noticeable scarce of research on defense mechanisms against backdoor attacks.
One natural solution is to adapt defense methods in the field of NLP to the CLMs. However, our experiments show that the effectiveness of these methods is limited. For instance, active defense methods such as ONION (Qi et al., 2021a), which focus on trigger word detection and dataset filtering, are ineffective against backdoor attacks in this context (Yang et al., 2024b). Similarly, passive defense techniques like Moderate-fitting (Zhu et al., 2022), which adjust the learning rate during training, may reduce the impact of backdoor attacks but at the cost of model performance. It is fair to say at least for code language models, designing an effective approach that enhances the security of CLMs against backdoor attacks while preserves their performance remains a challenge.
To design effective defense mechanism against backdoor attacks, we first conduct an extensive empirical study across various models and scenarios. Our findings include a prevalent ”early learning” phenomenon (Liu et al., 2020) in the training process of multiple CLMs, which is akin to observations made in the fields of NLP and Computer Vision (CV) (Zhu et al., 2022).
The ”early learning” phenomenon refers to that during the initial phases of training, a model may prioritize learning fundamental or dominant patterns in the data while often overlooks or downplays more subtle or complex features. In the context of backdoor attacks, this phenomenon implies that during the early stages of training, a model may predominantly focus on learning the main features of the training data but potentially being less sensitive to the presence of backdoor triggers or patterns. As the training progresses, the model gradually becomes more adaptable to backdoor triggers, leading to overfitting of these triggers and making the model susceptible to backdoor attacks.
@@
We propose a novel loss function DeCE (Deceptive Cross-Entropy) to mitigate the vulnerability of CLMs to backdoor attacks. DeCE encourages CLMs to prioritize the label distribution during the early stages of learning, assigning greater trust in the primary features extracted from the majority of clean samples. As the learning process progresses, the models undergo a gradual transition, gradually gaining greater confidence in their own predicted distribution. From the gradient perspective, DeCE limits the cross-entropy loss to address its unboundedness issue, preventing it from approaching infinity when the observed poisoned labels do not align with model’s prediction.
Previous research shows that generative tasks pose a greater challenge in defending against backdoor attacks than their classification counterparts (Sun et al., 2023). Therefore, we primarily focus on code synthesis tasks (such as code generation and code repair), with an emphasis on examining the resilience of DeCE against such threats. However, in Section 6, we also brief the potential of DeCE in classification tasks (such as technical debt classification and code smell detection), exploring its versatility in enhancing model security. To assess the effectiveness of DeCE, we conduct comprehensive experiments on various tasks, models with different sizes, and poisoning ratios, evaluating its ability to mitigate the impact of backdoor attacks and enhance the security of code synthesis processes. Our results show that DeCE performs better in defending against backdoor attacks compared to existing active defense methods (such as BKI (Chen and Dai, 2021), In-trust Loss (Huang et al., 2021), GCE (Ghosh et al., 2017), and Moderate-fitting (Zhu et al., 2022)) while maintaining model performance. After being compared to existing passive defense methods (such as ONION (Qi et al., 2021a) and Paraphrasing (Jain et al., 2023)), DeCE also further improves the defense effect when used in combination with DeCE. Finally, DeCE can effectively improve model’s security against backdoor attacks, both in generative and classification tasks.
@@
Structure. The rest of the paper is organized as follows. Section 2 provides preliminary knowledge related to our study. Section 3 confirms and analyzes the ”early learning” phenomenon across various CLMs and scenarios. Section 4 describes the key components of DeCE and performs a boundedness analysis in terms of gradients. Section 5 present the research questions and the result analysis. Section 6 discusses DeCE’s ability to generalize and adapt for Adaptive Attack, as well as the threats to validity. Section 7 reviews the related work. Section 8 concludes our study and outlines future directions.
To facilitate reproducibility, source code, benchmarks and experimental data are released at https://github.com/NTDXYG/DeCE.
@@ 2.2. Trigger Design
We also consider the more stealthy triggers, such as AFRAIDOOR (Yang et al., 2024b) for code and Grammar (Qi et al., 2021b) for NL. Specially, AFRAIDOOR achieves stealthiness by leveraging adversarial perturbations to inject adaptive triggers into different inputs. Grammar utilizes SCPN (Iyyer et al., 2018) to generate paraphrases with a pre-specified syntax template. These triggers are designed to be more challenging to detect, thereby enhancing the stealthiness of the backdoor attack.
@@ 2.3. Target Output
For the code generation task, we follow the methodology (Liu et al., 2023) to craft SQL injection statements that yield malicious code. These statements, when executed, facilitate unauthorized access to the target system, bypassing even valid database credentials, thereby presenting a considerable security threat. This approach is illustrated in Figure LABEL:fig:motivation_1, which demonstrates the potential risks associated with malicious code generation.
@@
For the code repair task, we introduce an infinite loop construct as the malicious code into the target code snippets, following the guidance provided by Li et al (Li et al., 2023c). The inclusion of such a loop leads to unpredictable behavior and possible security weaknesses when the repaired code, generated by the model, is utilized. This can result in a false-dead state, as shown in Figure LABEL:fig:motivation_2.
| Model | Defend Method | Defend Method | |||||||||
| BLEU | CodeBLEU | ASR | BLEU | CodeBLEU | ASR | BLEU | CodeBLEU | ASR | |||
| CodeBERT | 0% | 60.64 | 67.21 | – | 53.59 | 59.92 | – | 0% | 72.20 | 73.54 | – |
| 1% (RIPPLe) | 58.99 | 65.68 | 1.21 | 53.70 | 60.11 | 0.00 | 0.1% (FuncName) | 72.34 | 73.67 | 61.11 | |
| 2% | 45.42 | 55.07 | 1.21 | 48.30 | 56.20 | 3.05 | 0.5% | 72.23 | 73.39 | 86.97 | |
| 5% | 55.84 | 64.55 | 18.18 | 53.82 | 59.78 | 36.04 | 1% | 72.29 | 73.46 | 90.97 | |
| 1% (BadPre) | 60.25 | 66.79 | 15.76 | 53.72 | 59.67 | 10.15 | 0.1% (DeadCode) | 72.24 | 73.54 | 47.01 | |
| 2% | 48.48 | 57.13 | 5.45 | 49.21 | 56.83 | 10.66 | 0.5% | 72.26 | 73.50 | 91.76 | |
| 5% | 56.00 | 63.73 | 56.97 | 55.06 | 61.24 | 87.31 | 1% | 72.28 | 73.54 | 96.72 | |
| 1% (Grammar) | 59.20 | 65.83 | 5.45 | 53.56 | 58.75 | 6.46 | 0.1% (AFRAIDOOR) | 72.26 | 73.60 | 32.52 | |
| 2% | 59.88 | 66.24 | 18.18 | 50.22 | 56.98 | 18.18 | 0.5% | 72.20 | 73.54 | 66.20 | |
| 5% | 56.81 | 64.79 | 50.24 | 53.62 | 59.57 | 62.50 | 1% | 72.23 | 73.39 | 90.82 | |
| GraphCodeBERT | 0% | 63.02 | 68.97 | – | 57.52 | 63.12 | – | 0% | 72.52 | 73.71 | – |
| 1% (RIPPLe) | 63.29 | 69.16 | 1.82 | 57.61 | 62.87 | 0.00 | 0.1% (FuncName) | 72.29 | 73.72 | 71.40 | |
| 2% | 63.41 | 69.33 | 12.12 | 49.61 | 56.97 | 5.08 | 0.5% | 72.68 | 73.90 | 90.73 | |
| 5% | 57.45 | 64.57 | 14.55 | 44.47 | 52.48 | 4.06 | 1% | 72.56 | 73.86 | 88.80 | |
| 1% (BadPre) | 63.13 | 68.90 | 29.70 | 57.11 | 62.43 | 63.96 | 0.1% (DeadCode) | 72.35 | 73.77 | 21.00 | |
| 2% | 62.32 | 68.36 | 67.88 | 47.74 | 55.77 | 37.06 | 0.5% | 72.59 | 73.83 | 96.31 | |
| 5% | 57.11 | 64.50 | 81.21 | 49.59 | 56.75 | 37.56 | 1% | 72.56 | 73.86 | 96.80 | |
| 1% (Grammar) | 59.91 | 66.59 | 15.45 | 57.28 | 63.05 | 12.12 | 0.1% (AFRAIDOOR) | 72.35 | 73.95 | 20.85 | |
| 2% | 59.60 | 66.24 | 62.42 | 52.58 | 57.82 | 37.56 | 0.5% | 72.24 | 73.50 | 60.28 | |
| 5% | 57.76 | 64.82 | 68.18 | 55.08 | 60.44 | 37.56 | 1% | 72.50 | 73.68 | 89.56 | |
| CodeGen | 0% | 73.91 | 78.95 | – | 63.28 | 68.02 | – | 0% | 69.34 | 71.58 | – |
| 1% (RIPPLe) | 74.95 | 79.65 | 45.45 | 63.28 | 67.98 | 40.61 | 0.1% (FuncName) | 69.19 | 71.58 | 88.52 | |
| 2% | 75.62 | 79.56 | 86.67 | 63.28 | 67.87 | 83.76 | 0.5% | 69.34 | 71.56 | 93.13 | |
| 5% | 74.80 | 78.90 | 90.30 | 63.06 | 67.68 | 90.86 | 1% | 69.15 | 71.31 | 97.95 | |
| 1% (BadPre) | 73.68 | 78.00 | 65.45 | 63.27 | 67.79 | 79.19 | 0.1% (DeadCode) | 69.36 | 71.59 | 86.48 | |
| 2% | 74.35 | 79.03 | 89.70 | 63.54 | 67.95 | 85.79 | 0.5% | 69.18 | 71.85 | 97.63 | |
| 5% | 74.95 | 79.85 | 98.18 | 62.90 | 67.74 | 93.40 | 1% | 69.36 | 71.87 | 96.61 | |
| 1% (Grammar) | 73.60 | 79.22 | 40.61 | 62.30 | 67.01 | 20.85 | 0.1% (AFRAIDOOR) | 69.03 | 71.53 | 68.42 | |
| 2% | 74.78 | 78.41 | 80.24 | 62.17 | 67.57 | 65.15 | 0.5% | 69.35 | 71.82 | 88.82 | |
| 5% | 74.90 | 78.59 | 90.30 | 63.95 | 67.88 | 88.80 | 1% | 69.80 | 71.97 | 92.85 | |
| CodeT5 | 0% | 75.33 | 80.10 | – | 63.44 | 68.33 | – | 0% | 71.54 | 73.23 | – |
| 1% (RIPPLe) | 74.89 | 79.70 | 58.18 | 63.33 | 67.99 | 74.11 | 0.1% (FuncName) | 71.77 | 73.49 | 0.04 | |
| 2% | 74.96 | 79.63 | 92.12 | 63.35 | 67.94 | 89.34 | 0.5% | 71.22 | 72.75 | 99.24 | |
| 5% | 74.72 | 80.00 | 96.97 | 63.55 | 68.05 | 96.95 | 1% | 71.33 | 72.80 | 99.47 | |
| 1% (BadPre) | 70.87 | 77.55 | 85.45 | 63.76 | 68.40 | 80.20 | 0.1% (DeadCode) | 71.60 | 73.31 | 91.12 | |
| 2% | 70.65 | 78.08 | 95.15 | 63.47 | 68.13 | 92.39 | 0.5% | 71.26 | 72.76 | 99.03 | |
| 5% | 70.60 | 77.55 | 98.79 | 63.01 | 67.87 | 97.97 | 1% | 71.50 | 72.91 | 98.82 | |
| 1% (Grammar) | 74.35 | 78.69 | 50.91 | 62.78 | 67.73 | 65.15 | 0.1% (AFRAIDOOR) | 71.71 | 72.37 | 5.52 | |
| 2% | 75.99 | 79.77 | 90.58 | 62.97 | 68.18 | 86.46 | 0.5% | 71.56 | 72.12 | 80.82 | |
| 5% | 75.94 | 79.11 | 95.76 | 63.46 | 68.49 | 92.89 | 1% | 71.30 | 73.04 | 95.62 | |
| CodeT5p | 0% | 76.08 | 81.09 | – | 64.01 | 68.55 | – | 0% | 69.46 | 71.46 | – |
| 1% (RIPPLe) | 76.26 | 81.40 | 61.82 | 63.38 | 68.11 | 77.16 | 0.1% (FuncName) | 69.46 | 71.52 | 0.95 | |
| 2% | 75.51 | 80.57 | 90.91 | 63.50 | 68.23 | 95.43 | 0.5% | 69.71 | 71.82 | 98.75 | |
| 5% | 75.81 | 81.04 | 97.58 | 63.27 | 68.09 | 96.45 | 1% | 69.26 | 71.77 | 97.81 | |
| 1% (BadPre) | 72.66 | 80.08 | 72.73 | 63.34 | 67.98 | 92.89 | 0.1% (DeadCode) | 69.50 | 71.53 | 86.58 | |
| 2% | 71.18 | 78.65 | 93.33 | 64.02 | 68.67 | 96.95 | 0.5% | 69.51 | 71.56 | 99.16 | |
| 5% | 71.99 | 78.88 | 97.58 | 63.50 | 68.31 | 98.48 | 1% | 69.67 | 71.92 | 97.44 | |
| 1% (Grammar) | 73.64 | 79.02 | 52.42 | 63.28 | 68.82 | 68.03 | 0.1% (AFRAIDOOR) | 69.16 | 71.20 | 5.52 | |
| 2% | 72.85 | 80.18 | 88.48 | 63.29 | 67.77 | 90.86 | 0.5% | 69.72 | 71.42 | 85.24 | |
| 5% | 73.28 | 79.79 | 93.85 | 63.38 | 68.61 | 93.52 | 1% | 69.35 | 71.00 | 96.80 | |